• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name ransome

    Log entry time 00:00:04 on June16,2000

    Entry number 43885

    keyword=shift summary

    Shift start with beam still off
    16:10 Ed Folts calls and says termperature in hadron hut is 90 F.
    Plant engineering came, fixed relay contact on AC
    17:45 Hall closed up
    18:05 beam back with 2 uA for Compton checks
    18:10 Compton checks done -
    18:18 Beam back at 70 uA, start running!
    18:35 beam problems, go to low current for a minute
    18:40 back on




    Run Events Comment
    2101 1.3 M
    2102 1.3 M
    2103 1.3 M
    2104 1.3 M
    2105 1.3 M
    2106 1.3 M
    2107 1.3 M
    2108 1.3 M
    2109 1.3 M
    2110 1.3 M
    2111 1.3 M
    2112 1.3 M Very stable running. No beam trips in the last hour.